OpenStudy (anonymous):

(4^(12a )/ 5^(12a)) ^1/3 Rewrite the following expression without parentheses. Simplify your answer as much as possible, and assume that all variables are positive

zepdrix (zepdrix):

\[\Large\rm \left(\frac{4^{12a}}{5^{12a}}\right)^{1/3}\]Hey owl :O We need to simplify this?

zepdrix (zepdrix):

Recall that when we have something like this: \[\Large\rm (x^2)^3=x^{2\cdot3}=x^6\]Exponent and another exponent outside, we multiply the values.

OpenStudy (anonymous):

yeah

zepdrix (zepdrix):

To get rid of the brackets, you need to distribute the 1/3 to both the exponential in the top AND bottom.

zepdrix (zepdrix):

\[\Large\rm \left(\frac{4^{12a}}{5^{12a}}\right)^{1/3}=\frac{\left(4^{12a}\right)^{1/3}}{\left(5^{12a}\right)^{1/3}}\]

zepdrix (zepdrix):

And then multiply the exponents, yah? :o What's 12 times 1/3?

OpenStudy (anonymous):

4

zepdrix (zepdrix):

\[\Large\rm =\frac{4^{4a}}{5^{4a}}\]Mmmm ok good good good. Applying that exponent rule to each the top and bottom gives us something like this, yes?

OpenStudy (anonymous):

yup

zepdrix (zepdrix):

Since the `bases` of these exponentials are different, there isn't much else we can do with them. We would stop there :D

OpenStudy (anonymous):

ok thank you very much!